OpenStudy (anonymous):

solve by factoring 25x^2 -20x +4 =0

OpenStudy (anonymous):

looks like it might be a perfect square

OpenStudy (dls):

(5x-2)^2

OpenStudy (anonymous):

Write the quadratic equation in standard form. Divide both sides by 25: x^2-(4 x)/5+4/25 = 0 Factor the left hand side. Write the left hand side as a square: (x-2/5)^2 = 0 Eliminate the exponent on the left hand side. Take the square root of both sides: x-2/5 = 0 Solve for x. Add 2/5 to both sides: Answer: | | x = 2/5

OpenStudy (anonymous):

thank you!